OpenStudy (anonymous):

The seinfield's family room measures 23 feet by 25 feet. They are covering the floor with square tiles that measure 1 foot on a side and costs $0.76 each. How much will they spend on the tile?

OpenStudy (anonymous):

First we need to figure out how many titles fit into a room. We can do this by dividing the total square footage of the floor by the square foot of a single title. \[\frac{ 23 \times 25 }{ 1 \times 1 } = \frac{ 575 }{ 1 } = 575\]This means that the floor has enough room for 575 titles. This means to find the total price of 575 titles we just multiply the number of the titles by the cost; which is $0.76 in this case. \[575 \times 0.76 = 437\]This means that 575 titles will cost you $437.

OpenStudy (anonymous):

Thanks